# Revisiting Node Affinity Prediction in Temporal Graphs

### v1

@@ 3.2 NAViS: Node Affinity Prediction with a Global Virtual State
where ${\mathbf{x}},{\mathbf{h}}_{i-1},{\mathbf{h}}_{i},{\mathbf{g}},{\mathbf{s}}\in\mathbb{R}^{d}$ are the previous affinity vector, previous node state, updated node state, virtual global vector, and predicted affinity vector, respectively. The parameters $W_{xh},W_{hh},W_{xs},W_{hs},W_{gs}\in\mathbb{R}^{d\times d}$ are learnable weights, ${\mathbf{b}}_{h},{\mathbf{b}}_{s}\in\mathbb{R}^{d}$ are learnable biases, and $\sigma$ is the sigmoid function, forcing ${\mathbf{z}}_{h},{\mathbf{z}}_{s}$ to be in $[0,1]$, to maintain conceptual similarity with $\alpha$ in Equation˜6. In Figure˜2 we provide a detailed scheme of NAViS.
@@
NAViS for large-scale graphs. For graphs with $N$ nodes, the number of weights of NAViS scales as $O(N^{2})$, which is prohibitive at scale. To make NAViS practical on large graphs, we introduce a sparsified affinity prediction pipeline. Specifically, for each node, we retain only the top-$a$ entries (by magnitude) of its previous affinity vector, sort them by magnitude, and maintain an index map $\pi$ from original to compressed coordinates. We use $\pi$ to map all other inputs consumed by NAViS, and during training we apply the same mapping to the ground-truth future affinity vector before computing the loss, so that optimization is performed on the compressed vectors. This restricts per-node computation to the subset of neighbors that previously exhibited high affinity, reducing time and memory to $O(1)$ for a constant $a$, while avoiding the quadratic blow-up in parameters. Empirically—and consistent with prior observations that affinity vectors are typically extremely sparse—this truncation rarely discards non-negligible mass even on very large graphs (Huang et al., 2023).
@@ 5 Related Work
Expressivity in Temporal Graphs. The dominant view of expressivity in graph learning is measured by the Weisfeiler-Lehman (WL) test’s ability to distinguish non-isomorphic graphs (Gilmer et al., 2017; Xu et al., 2019). This concept extends to temporal graphs, where models are evaluated on their capacity to differentiate evolving graph structures (Kazemi et al., 2020; ENNADIR et al., 2025). We diverge from this perspective by focusing on functional expressivity: a model’s ability to represent specific mathematical operations. While prior models aim to capture complex graph topology, they often cannot represent a simple Moving Average, a critical function for affinity prediction.

### v2

@@ 3.2 NAViS: Node Affinity Prediction with a Global Virtual State
where ${\mathbf{x}},{\mathbf{h}}_{i-1},{\mathbf{h}}_{i},{\mathbf{g}},{\mathbf{s}}\in\mathbb{R}^{d}$ are the previous affinity vector, previous node state, updated node state, virtual global vector, and predicted affinity vector, respectively. The parameters $W_{xh},W_{hh},W_{xs},W_{hs},W_{gs}\in\mathbb{R}^{1\times d}$ are learnable weights, ${\mathbf{b}}_{h},{\mathbf{b}}_{s}\in\mathbb{R}^{d}$ are learnable biases, and $\sigma$ is the sigmoid function, forcing ${\mathbf{z}}_{h},{\mathbf{z}}_{s}$ to be in $[0,1]$, to maintain conceptual similarity with $\alpha$ in Equation˜6. In Figure˜2 we provide a detailed scheme of NAViS.
@@
NAViS for large-scale graphs. For graphs with $N$ nodes, the number of learnable parameters in NAViS scales as $O(N)$, which can be prohibitive for graphs with millions of nodes. To make NAViS practical at this scale, we introduce a sparsified affinity prediction pipeline. Specifically, for each node we retain only the entries corresponding to candidate target nodes. In real-world settings, such as streaming services where users and movies are nodes and interactions are edges, we are interested in each user’s affinity to movies rather than to other users. In practice, this substantially reduces the parameter count of NAViS. For example, on the tgbn-token dataset (Shamsi et al., 2022), which records user–token interactions, NAViS requires about (5,000) parameters, while the graph contains over (60,000) nodes.
@@ 5 Related Work
Expressivity in Temporal Graphs. The dominant view of expressivity in graph learning is measured by the Weisfeiler-Lehman (WL) test’s ability to distinguish non-isomorphic graphs (Gilmer et al., 2017; Xu et al., 2019). This concept extends to temporal graphs, with temporal-WL (Souza et al., 2022) for CTDG and supra Laplacian WL (Galron et al., 2025) for DTDG, where models are evaluated on their capacity to differentiate evolving graph structures (Kazemi et al., 2020; ENNADIR et al., 2025). We diverge from this perspective by focusing on functional expressivity: a model’s ability to represent specific mathematical operations. While prior models aim to capture complex graph topology, they often cannot represent a simple Moving Average, a critical function for affinity prediction.